How do you know what formula to use for probability?
For example for the problem:
A Company had manufactures a batch of 400 backup power supply units for computers, 8 of them are defective. If 3 of the units are randomly selected for testing, what is the probability that the entire batch will be accepted?
What formula would i use to solve this? Thank you!

Answer by edjones(8007)

8/400=1/50 defective
49/50 not defective.
(49/50)^3=.9412 probability none of the 3 will be found defective.
.
Ed